Statistics and Probability Question. Please show and explain all calculations in detail so that we can learn. Thank you!Problem 5. Market data from customers who buy Surface Pro 7 or Surface Pro 6 computers show that a randomly selected customer
Answer #1

a) since here the buying of the surface pro is treated as a success with a probability (p) of 0.6 and buying surface pro 6 as a failure (q) with probability 0.4 and p=1-q

so here it follows a binomial distribution, i.e, \binom{n}{x}p^xq^{n-x} , here n=100, p=0.6, q=0.4

b)as we know the mean of a binomial distribution as np and standard deviation as \sqrt npq

therefore, mean =100*0.6=60

standard deviation= \sqrt (100*0.6*0.4) = 4.89

c) here we have to find

P(X\geq mean+0.5sd)=P(X\geq 63)=\sum_{i=63}^{100}\binom{n}{i}p^iq^{n-i}P(X\geq mean+0.5sd)=P(X\geq 63)=\sum_{i=63}^{100}\binom{n}{i}p^iq^{n-i}=0.239

d)here we have to calculate the probability of 70 and above independent customers getting what they want

as we can see that the minimum no. of smart pro 6 is 30 so to satisfy the 100 customers we calculate the probability as follows

P(X\geq 70)=\sum_{i=63}^{100}\binom{n}{i}p^iq^{n-i}=0.14775